Determine whether the graph of y = x^2 − 6x + 3 has a maximum or minimum point, then find the maximum or minimum value.
Answer by stanbon(75887)

Determine whether the graph of y = x^2 − 6x + 3
has a maximum or minimum point, then find the maximum or minimum.
--------
Since the coefficient of x^2 is positive,
y will increase as x gets larger and larger
both positively and negatively.
So the graph has a minimum point.
-------
The minimum occurs where x = -b/(2a) = 6/(2*1) = 3
---
The minimum y-value = f(3) = 3^2-6*3+3 = 9-18+3 = -6
-------------------
Cheers,
Stan H.